OpenStudy (anonymous):

What is the slope of the line passing through (0,3) and (4,5)? A. 0.5 B. 2 C. -2 D. -0.5

OpenStudy (hoblos):

(5-3)/4 = 2/4 =0.5

OpenStudy (hoblos):

A

OpenStudy (anonymous):

To find the slope, you can calculate it using the equation rise / run. What's the rise (change in y value) and what's the run (change in x value)?

OpenStudy (anonymous):

step by step explaining please :)

OpenStudy (anonymous):

In other words: Step 1: What's the difference between the first y coordinate and the second y coordinate? (difference = subtract!) Step 2: Do the same thing with the x coordinates. Step 3: Divide the y difference by the x difference.